• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 08:02:06 on December20,2003

    Entry number 114670

    keyword=Owl shift summary

    Ramped Septa by step of 50 A ( 50 A on one and 50 A on the other...) until 200 A, the magnetic fields do cancel limiting the beam steering.

    Took 3 runs :
    1436 Thick C12
    1437 Thin C12
    1438 Waterfall target

    (look at Tim Holmstrom Halog entries )

    Rest of the shift as down time, with Scott Spiegel trying to fix the left arm